# Current density distribution for the quantum Hall effect

### v2

@@ III Current density distribution for the QHE
The current density distribution for the case of a quantized Hall resistance became the subject of controversial discussions [27, 28] shortly after the discovery of the QHE, which are still continuing. The common non-interacting Landauer-Büttiker picture assumes current flow within compressible one-dimensional edge channels [6]. Taking into account the Coulomb interaction of the carriers, Chklovskii et al. predicted the formation of ICSs [15] and then realized, calculating the ballistic current through a narrow gate defined channel, that for a quantized Hall resistance the current flows inside the ICSs [16]. Gerhardts and collaborators used a local version of Ohm’s law assuming that the imposed current density, $j_{\text{impo}}(y)$, is proportional to the gradient of the chemical potential, $d\mu/dy$, which they calculated based on the screening theory [19]. Since $d\mu/dy=0$ within compressible regions, they confirmed that the current flows inside the ICSs. However, by assuming $j_{\text{impo}}\propto|d\mu/dy|$ in their calculations, they completely neglected the persistent current density, $j_{\text{pers}}$, which led them to the prediction that the imposed current flows unidirectionally through all existing ICSs, i.e., in the same direction on both sides of the Hall bar [19]. Another approach predicting $j_{\text{impo}}$ uses the continuity equation and finds $j_{\text{impo}}\neq 0$ inside compressible regions [23]. We find this approach questionable, mainly because it is restricted to $I_{\text{impo}}\sim 0$ (linear response) and, hence, fails to account for the non-linear nature of the QHE, which matters for reasonable values of $V_{\text{H}}$.
@@
In the present article, we acknowledge in agreement with Refs. [16, 19, 4, 29] that the current flows inside the ICSs. However, instead of focusing on the gradient of the chemical potential, which was used in Ref. [19] to predict $j_{\text{impo}}$, we go back to the original approach by Chklovskii et al. [16] and consider the complete current density $j=j_{\text{impo}}+j_{\text{pers}}$, which is carried by the fully occupied LLs inside the ICSs. We emphasize, that inside the unscreened ICSs the electrons of a LL are driven by the local electric field, which is identical to the gradient of the LL energy, $\vec{E}=-\vec{\nabla}\varepsilon/e$ and differs from the gradient of the chemical potential.
@@
In Fig. 2, we qualitatively sketch our prediction for $\nu=1$ based on Eq. (2) and the screening theory [12]. The four panels show for four different characteristic situations $\varepsilon(y)\equiv\varepsilon_{1}(y)$ (solid lines), $\mu(y)$ (dotted red lines) and the related $j(y)$ (lower panels) across a Hall bar. ICSs are indicated by a dark gray background, compressible regions by light gray. Only within the unscreened ICSs the LL energy, $\varepsilon(y)$, varies and $j\neq 0$. In Figs. 2a and 2b, we consider the equilibrium with $I_{\text{impo}}=0$, while in Figs. 2c and 2d we added a finite $I_{\text{impo}}\neq 0$. In Figs. 2a and 2c, we depict the case of two ICSs near the edges of the Hall bar, while in Figs. 2b and 2d, we sketch the case of a single bulk ICS. The screening theory predicts these two cases to occur at the two ends of the plateau of a quantized Hall resistance as indicated in the insets above Figs. 2a and 2b [12].
@@
Because, the persistent current flows in a closed loop, its density fulfills $\int_{-b}^{b}dyj_{\text{pers}}=0$. Hence, we can define $I_{\text{pers}}$ by dividing this integral into two regions: $I_{\text{pers}}\equiv-\int_{-b}^{y(j=0)}dyj_{\text{pers}}=\int_{y(j=0)}^{b}dyj_{\text{pers}}$, which has the identical absolute value on both sides of the Hall bar. In contrast to $I_{\text{pers}}$, the imposed current enters the Hall bar and leaves is through the source and drain contacts. Therefore, $I_{\text{impo}}$ corresponds to the integral of the complete current density over the full width of the Hall bar, $I_{\text{impo}}=\int_{-b}^{b}dyj=\int_{-b}^{b}dyj_{\text{impo}}$.
@@
$I_{\text{impo}}$ generates the Hall voltage between the two edges, $|V_{\text{H}}|=|\mu(b)-\mu(-b)|/e=h/e^{2}/I_{\text{impo}}$ (for $\nu=1$), cf. Figs. 2c and 2d. In our sketches we chose the chemical potential at the drain contact as reference, $\mu(-b)=\mu_{0}$. In equilibrium, i.e., for $I_{\text{impo}}=0$, the chemical potential is constant, $\mu(y)=\mu_{0}$. For $I_{\text{impo}}\neq 0$, $\mu(y)\geq\mu_{0}$ increases monotonously between the edges, while it always remains constant within the compressible regions. An increase of $\mu(y)$ corresponds to an accumulation of a negative charge. Therefore, $\varepsilon(y)$ increases accordingly (compare the upper panels for $I_{\text{impo}}=0$ with the lower panels for $I_{\text{impo}}\neq 0$ in Fig. 2). Importantly, the increase of $\varepsilon(y)$ yields a decrease of $j$ for $b<y\lesssim 0$ (the low potential side of the Hall bar) and an increase of $j$ for $0\lesssim y<b$ (the high potential side). Because $\int_{-b}^{b}dyj_{\text{pers}}=0$, the decrease of $j$ on the low potential side corresponds to a decrease of $I_{\text{pers}}$. The increase of $j$ on the high potential side then points to a combination of the established decrease of $I_{\text{pers}}$ together with $I_{\text{impo}}$ flowing there.
@@
Within our linear approximation, see Sec. II, for $eV_{\text{H}}\ll 2\Delta E$ we find an opposite direction of the current flow on the two sides of the Hall bar, i.e., a chiral current flow. The case $eV_{\text{H}}\sim 2\Delta E$, which is clearly beyond the scope of our approximation, would result in a break down of the ICSs. However, as long as an ICS exists on the low potential side, we expect $d\varepsilon/dy$ to remain $\leq 0$ there, supported by numerical calculations [24]. We conclude: whenever the Hall resistance is quantized (ICSs exist), the current flow is chiral.
@@
Figure 3: Illustration of geometry in (b) of the ICS corresponding to the non-equilibrium situation with $I_{\text{impo}}\neq 0$ of Fig. 2(b); (a) and (c) are energy diagrams of the ICS along the upper and lower edges, repsecively. States contributing to $j_{\text{pers}}$ are indicated in blue and those of $j_{\text{impo}}$ in red. The persistent current describes a closed path inside the Hall bar as all contributing states have energies below the chemical potentials of both leads, $E<\mu_{\text{S}},\mu_{\text{D}}$. The imposed current is carried by states with energies $\mu_{\text{S}}\geq E>\mu_{\text{D}}$. Its electrons enter the Hall bar from the source contact, move along its high potential (lower) edge and leave it into empty states of the drain contact.

### v3

@@ III Current density distribution for the QHE
The current density distribution for the case of a quantized Hall resistance became the subject of controversial discussions shortly after the discovery of the QHE, which are still continuing Laughlin (1981); Halperin (1982); Büttiker (1988); Chklovskii et al. (1993); Wexler and Thouless (1994); Geller and Vignale (1994); Komiyama and Hirai (1996); Güven and Gerhardts (2003); Gerhardts (2020). We provide only a few historical steps relevant for our discussion without claiming completeness. The common non-interacting Landauer-Büttiker picture assumes current flow within compressible one-dimensional edge channels Büttiker (1988). Taking into account the Coulomb interaction of the carriers, Chklovskii et al. predicted the formation of ICSs Chklovskii et al. (1992) and then realized, calculating the ballistic current through a narrow gate defined channel, that for a quantized Hall resistance the current flows inside the ICSs Chklovskii et al. (1993). Gerhardts and collaborators used a local version of Ohm’s law assuming that the imposed current density, $j_{\text{impo}}(y)$, is proportional to the gradient of the chemical potential, $d\mu/dy$, which they calculated based on the screening theory Güven and Gerhardts (2003). Since $d\mu/dy=0$ within compressible regions, they confirmed that the current flows inside the ICSs. However, by assuming $j_{\text{impo}}\propto|d\mu/dy|$ in their calculations, they completely neglected the persistent current density, $j_{\text{pers}}$, which led them to the prediction that the imposed current flows unidirectionally through all existing ICSs, i.e., in the same direction on both sides of the Hall bar Güven and Gerhardts (2003). Another approach predicting $j_{\text{impo}}$ uses the continuity equation and finds $j_{\text{impo}}\neq 0$ inside compressible regions Armagnat and Waintal (2020). We find this result questionable, mainly because it is based on calculating the linear response conductivity in the limit $I_{\text{impo}}\sim 0$ by subtracting $j_{\text{pers}}$ from $j_{\text{pers}}+j_{\text{impo}}$ while $j_{\text{pers}}\gg j_{\text{impo}}$. However, for a finite $I_{\text{impo}}$ current flow inside the ICSs becomes evident, once we accept that the gradient of the energy of a LL results in a drift velocity of the carriers within that LL, i.e., the local electric field drives current.
@@
In the present article, we acknowledge in agreement with Refs. Chklovskii et al., 1993; Güven and Gerhardts, 2003; Gerhardts, 2008, 2020 that the current flows inside the ICSs. However, instead of focusing on the gradient of the chemical potential, which was used in Ref. Güven and Gerhardts, 2003 to predict $j_{\text{impo}}$, we go back to the original approach by Chklovskii et al. Chklovskii et al. (1993) and consider the complete current density $j=j_{\text{impo}}+j_{\text{pers}}$, which is carried by the fully occupied LLs inside the ICSs. We emphasize, that inside the unscreened ICSs the electrons of a LL are driven by the local electric field, which is identical to the gradient of the LL energy, $\vec{E}=-\vec{\nabla}\varepsilon/e$ and differs from the gradient of the chemical potential.
@@
In Fig. 2, we qualitatively sketch our prediction for $\nu=1$ based on Eq. (2) and the screening theory Siddiki and Gerhardts (2004). The four panels show for four different characteristic situations $\varepsilon(y)\equiv\varepsilon_{1}(y)$ (solid lines), $\mu(y)$ (dotted red lines) and the related $j(y)$ (lower panels) across a Hall bar. ICSs are indicated by a dark gray background, compressible regions by light gray. Only within the unscreened ICSs the LL energy, $\varepsilon(y)$, varies and $j\neq 0$. In Figs. 2a and 2b, we consider the equilibrium with $I_{\text{impo}}=0$, while in Figs. 2c and 2d we added a finite $I_{\text{impo}}\neq 0$. In Figs. 2a and 2c, we depict the case of two ICSs near the edges of the Hall bar, while in Figs. 2b and 2d, we sketch the case of a single bulk ICS. The screening theory predicts these two cases to occur at the two ends of the plateau of a quantized Hall resistance as indicated in the insets above Figs. 2a and 2b Siddiki and Gerhardts (2004).
@@
In the contacts, the confinement potential is screened, if they are compressible, and no persistent current flows. Hence, in order to comply with Kirchhoff’s junction rule (or particle conservation), $I_{\text{pers}}$ has to form a closed loop inside the Hall bar. Clearly, its density fulfills $\int_{-b}^{b}dyj_{\text{pers}}=0$. Hence, we can define $I_{\text{pers}}$ by dividing this integral into two regions: $I_{\text{pers}}\equiv-\int_{-b}^{y(j=0)}dyj_{\text{pers}}=\int_{y(j=0)}^{b}dyj_{\text{pers}}$, which has the identical absolute value on both sides of the Hall bar. In contrast to $I_{\text{pers}}$, the imposed current enters the Hall bar and leaves it through the source and drain contacts. Therefore, $I_{\text{impo}}$ corresponds to the integral of the complete current density over the full width of the Hall bar, $I_{\text{impo}}=\int_{-b}^{b}dyj=\int_{-b}^{b}dyj_{\text{impo}}=|V_{\text{H}}|\nu e^{2}/h$, where for integer filling factor $\nu=1,2,3,\dots$, the term $h/(\nu e^{2})$ is the quantized Hall resistance.
@@
The current density $j_{\text{impo}}$ generates the Hall potential between the two edges, which corresponds to the change of the chemical potential according to $V_{\text{H}}(y)=\left[\mu(y)-\mu_{0}\right]/e$, cf. Figs. 2c and 2d. The quantity $V_{\text{H}}(y)$ is measured in the scanning probe experiment in Ref. Weitz et al., 2000, if $\mu_{0}$ is the reference ground potential. In our sketches we chose the chemical potential at the drain contact as this reference, $\mu(-b)=\mu_{0}$. In equilibrium, i.e., for $I_{\text{impo}}=0$, the chemical potential is constant, $\mu(y)=\mu_{0}$. For $I_{\text{impo}}\neq 0$, $\mu(y)\geq\mu_{0}$ increases monotonously between the edges, while it always remains constant within the compressible regions. An increase of $\mu(y)$ corresponds to an accumulation of a negative charge. Therefore, $\varepsilon(y)$ increases accordingly (compare the upper panels for $I_{\text{impo}}=0$ with the lower panels for $I_{\text{impo}}\neq 0$ in Fig. 2). Importantly, the increase of $\varepsilon(y)$ yields a decrease of $j$ for $b<y\lesssim 0$ (the low potential side of the Hall bar) and an increase of $j$ for $0\lesssim y<b$ (the high potential side). Because $\int_{-b}^{b}dyj_{\text{pers}}=0$, the decrease of $j$ on the low potential side corresponds to a decrease of $I_{\text{pers}}$. The increase of $j$ on the high potential side then points to a combination of the established decrease of $I_{\text{pers}}$ together with $I_{\text{impo}}$ flowing there.
@@
Within our linear approximation, see Sec. II, for $eV_{\text{H}}\ll 2\Delta E_{\nu}$ we find an opposite direction of the current flow on the two sides of the Hall bar, i.e., a chiral current flow. The case $eV_{\text{H}}\sim 2\Delta E_{\nu}$, which is clearly beyond the scope of our approximation, would result in a break down of the ICSs. However, as long as an ICS exists on the low potential side, we expect $d\varepsilon/dy$ to remain $\leq 0$ there, supported by numerical calculations Panos et al. (2014). We conclude: whenever the Hall resistance is quantized (ICSs exist), the current flow is chiral.
@@
Figure 3: Illustration of geometry in (b) of the ICS corresponding to the non-equilibrium situation with $I_{\text{impo}}\neq 0$ of Fig. 2(c); (a) and (c) are energy diagrams of the ICS along the upper and lower edges, repsecively. States contributing to $j_{\text{pers}}$ are indicated in blue and those of $j_{\text{impo}}$ in red. The persistent current describes a closed path inside the Hall bar as all contributing states have energies below the chemical potentials of both leads, $E<\mu_{\text{S}},\mu_{\text{D}}$. The imposed current is carried by states with energies $\mu_{\text{S}}\geq E>\mu_{\text{D}}$. Its electrons enter the Hall bar from the source contact, move along its high potential (lower) edge and leave it into empty states of the drain contact.
@@ IV Summary
Data availability statement
This article contains no experimental or theoretical data. Therefore, no data can be made publicly available.